[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Nov 17, 2015
Matthew Sardi - NOAA Federal
matthew.sardi at noaa.gov
Tue Nov 17 23:46:27 UTC 2015
FYI, we have an NCF trouble-ticket open on the AWIPS text database/EDEX
issues Chris mentioned.
If there are any AWIPS issues overnight, please call the NCF and mention
the open ticket:
NCF TT #: 713235
NCF POC: Mary
NHC/Technology & Science Branch
On Tue, Nov 17, 2015 at 6:22 PM, Christopher Mello - NOAA Federal <
christopher.mello at noaa.gov> wrote:
> One of the processes on the dx3 did not come back up correctly. This
> should not effect operations since the other EDEX servers can handle the
> work load.
> That being said, we had some strange issues with the text-database. Some
> of the products did not get stored as expected after they were sent out.
> The system appears stable at 6:00 PM...
> But we will address the issues Wednesday morning. My cell number is on the
> sheet if there is a problem overnight sending out any products...
> On Tue, Nov 17, 2015 at 7:47 AM, Christopher Mello - NOAA Federal <
> christopher.mello at noaa.gov> wrote:
>> DX AWIPS2 Hardware Refresh on Tuesday (TODAY). It will not effect
>> The hurricane hot line appears to be working normally.
>> HSU reported the it took longer than normal to produce some Atlantic ATCF
>> graphics. Monica or Craig will investigate.
>> Chris fixed the SREF model ingest procedure in GFE.
>> 12Z SFC analysis failed to reach OPC. Andy experienced this several times
>> over the past week. Re-sending using trop.launch fixed the problem each
>> time. TSB will investigate.
>> Other Issues:
>> The VM server (nhc-lx-kvm01) that hosts NHC's backup websites will be
>> rebooted Thursday to resolve a memory issues.
>> Chris Mello
>> National Hurricane Center
> Chris Mello
> National Hurricane Center
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Ncep.nhc.nco_contacts